# The Fellowship of the LLMs: Multi-Agent Workflows for Synthetic Preference Optimization Dataset Generation

### v3

@@ Abstract
This paper presents synthetic Preference Optimization (PO) datasets generated using multi-agent workflows and evaluates the effectiveness and potential of these workflows in the dataset generation process. PO dataset generation requires two modules: (1) response evaluation, and (2) response generation. In the response evaluation module, the responses from Large Language Models (LLMs) are evaluated and ranked - a task typically carried out by human annotators that we automate using LLMs. We assess the response evaluation module in a 2 step process. In step 1, we assess LLMs as evaluators using three distinct prompting strategies. In step 2, we apply the winning prompting strategy to compare the performance of LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate. In each step, we use inter-rater agreement using Cohen’s Kappa between human annotators and LLMs. For the response generation module, we compare different configurations for the LLM Feedback Loop using the identified LLM evaluator configuration. We use the win rate (the fraction of times a generation framework is selected as the best by an LLM evaluator) to determine the best multi-agent configuration for generation. After identifying the best configurations for both modules, we use models from the GPT, Gemma, and Llama families to generate our PO datasets using the above pipeline. We generate two types of PO datasets, one to improve the generation capabilities of individual LLM and the other to improve the multi-agent workflow. Our evaluation shows that GPT-4o-as-a-Judge is more consistent across datasets when the candidate responses do not include responses from the GPT family. Additionally, we find that the LLM Feedback Loop, with Llama as the generator and Gemma as the reviewer, achieves a notable 71.8% and 73.8% win rate over single-agent Llama and Gemma, respectively.
@@ 1 Introduction
Large Language Models (LLMs) demonstrate a range of Natural Language Processing (NLP) capabilities, including text generation, question answering, and language understanding. However, LLMs can sometimes deviate from user instructions and exhibit unintended behaviors (Tamkin et al. 2021). To mitigate this problem and align the LLM outputs more closely with human preferences, techniques like Reinforcement Learning from Human Feedback (RLHF) are used, which involves fine-tuning LLMs using the reward signal from human preferences (Christiano et al. 2017). Improved methods like Direct Preference Optimization (DPO) (Rafailov et al. 2024) eliminate the need for fitting the reward model and are more stable and performant. In DPO, the preference optimization dataset requires a pair of accepted and rejected responses for each prompt. The accepted response is one that better aligns with the desired human preferences. Other techniques like Kahneman-Tversky Optimization (KTO) (Ethayarajh et al. 2024) require each response to indicate whether it is good or bad (i.e., as a binary classification task) instead of pairwise preferences.
@@
In the process of constructing the dataset of human preferences, the evaluation and ranking of the outputs generated by LLMs are typically done by human annotators, who assess these outputs based on various criteria such as instruction following, helpfulness, relevance, accuracy, depth, and creativity. The PO dataset generation process is divided into two modules: response evaluation and response generation. The response evaluation module involves assessing and ranking responses generated by LLMs, while the response generation module focuses on creating responses that align with the identified preferences. This manual process, while effective, is labor-intensive, time-consuming, inconsistent, and subject to human biases. In this work, we thus ask the question, “Can we use LLM agents to automate and improve response evaluation and generation for constructing preference optimization (PO) datasets?”.
@@
For the response evaluation step, we leverage LLMs as evaluators and compare several configurations including LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate to pick the best evaluation strategy. Previously, single agents have been used to generate the responses for PO datasets. However, we use a multi-agent framework for response generation, which allows us to generate more refined and higher-quality responses. The multi-agent approach uses the collaboration between multiple LLMs, where one agent can provide suggestions for improvements, and the other can revise the response based on the feedback. This iterative process leads to a thorough refinement of the generated content, ensuring that the final output better aligns with human preferences and expectations.
@@
In this paper, we present multiple DPO and KTO datasets. Our focus is on generating two separate categories of datasets: one aimed at improving the performance of individual LLMs and the other to enhance the effectiveness of multi-agent workflows. The primary aim of the first dataset is to enhance the performance and capabilities of individual LLMs by providing high-quality PO training data that better aligns with human judgment and expectations. The goal of the second category of the dataset is to improve the multi-agent frameworks of LLM Feedback Loop generation approach, enabling better feedback provision, response refinement, and decision-making. Our contributions can be summarized as follows:
@@
    We generate synthetic PO datasets for single-agent improvement by combining the best configuration for the evaluation and generation module. We also generate PO datasets for multi-agent improvement.
@@ 2.1 Preference Optimization
Preference Optimization has emerged as a pivotal technique for aligning model outputs with human preferences. Rafailov et al. 2024 introduce DPO, a method that simplifies solving the standard RLHF (Reinforcement Learning from Human Feedback) problem by converting it into a classification task, enabling the extraction of the optimal policy in a straightforward way. Hong, Lee, and Thorne 2024 introduce ORPO algorithm that combines the traditional supervised fine-tuning and preference alignment stages into a single process. The dataset for DPO and ORPO require annotated preference pairs, where each pair consists of two model outputs labeled according to which one better aligns with human preferences. Ethayarajh et al. 2024 introduce KTO, a cost-effective approach to align Large Language Models (LLMs) with human feedback, improving performance without the need for preference pairs. Argilla Distilabel (Álvaro Bartolomé Del Canto et al. 2024) uses LLM to judge between the responses of two models to create synthetic PO datasets. The datasets are available on Hugging Face [^1] https://huggingface.co/argilla. To our knowledge, no one has yet explored the use of Multi-Agent workflows for the generation of PO datasets. However, multi-agent frameworks have been utilized for other tasks which we discuss below.
@@ 2.2 Agent Frameworks
Recently, there has been a growing interest in using LLM multi-agent frameworks for different tasks. Zheng et al. 2023a presents an evaluation of LLM-as-a-Judge on the MT-Bench (Zheng et al. 2023b) and Chatbot Arena (Li et al. 2024). Their results reveal that strong LLM judges like GPT-4 can match both controlled and crowd-sourced human preferences well, achieving over 80% agreement, the same level of agreement between humans. Additionally, they evaluate several variants of Llama and Vicuna on the dataset. They study the limitations of LLM-as-a-judge, including position, verbosity, and self-enhancement biases, as well as limited reasoning ability. Verga et al. 2024 explore the use of LLMs-as-a-Jury. Their approach, a Panel of LLM evaluators (PoLL), composed of a larger number of smaller models outperforms a single large judge. They also show that the PoLL approach exhibits less intra-model bias as compared to LLM-as-a-Judge. They use Command-R, GPT, Claude-3, and Mistral families for their study. Additionally, they compare two prompting strategies: (1) reference-based scoring where they provide the LLM with a reference answer, and (2) candidate answer and pair-wise scoring where they ask the LLM to pick the better response from the candidate responses. PoLL outperforms single-agents on KILT (Petroni et al. 2021) and Chatbot Arena.
@@
Liang et al.2024 introduce Multi-Agent Debate (MAD) to encourage divergent thinking in LLMs. They mitigate the Degeneration-of-Thought (DoT) problem, which is that once the LLM has established confidence in its solutions, it is unable to generate novel thoughts. In their approach, the affirmative LLM and the negative LLM debate on the answer while the LLM judge evaluates both arguments after each round of debate. They evaluate the approach on the Commonsense Machine Translation Dataset (Chinese to English) (He et al. 2020) and their Counter-Intuitive Arithmetic Reasoning (CIAR) dataset. MAD was able to achieve a 37% accuracy on the CIAR dataset using GPT-3.5-Turbo which outperforms Chain-of-Thought, Self-Consistency, and Self-Reflection prompting. They also show that using the MAD approach decreases bias and increases response diversity. Du et al. 2023 evaluates a different variant of multi-agent debate where multiple models generate their own responses, and each model receives the opinions of the other models, then updates its response if necessary. This is done for multiple rounds. Du et al. 2023 evaluates the approach on the following tasks: Biography generation, MMLU, Chess move validity, Arithmetic, Grade school math, and Chess move optimality. Their approach using ChatGPT and Bard outperforms single-agent on all the tasks. To evaluate LLM responses Chan et al. 2023 presents another variant of multi-agent debate. Their architecture involves assigning agents different roles such as General Public, Critic, Psychologist, News Author, and Scientist. They used ChatGPT and GPT-4 for their evaluation on FairEval (Wang et al. 2023a) dataset and achieved a Cohen’s Kappa score of 0.40 using LLM Debate, 0.03 more than the single agent.
@@ 3.1 Experimental Setup
In this study, we perform experiments on the three categories of models given in Table 1. For the evaluation module, we evaluate single agents and multi-agent frameworks on four datasets, Alpaca Eval (Li et al. 2023), FairEval (Wang et al. 2023a), PandaLM-Eval (Wang et al. 2024, 2023b) and MT-Bench (Zheng et al. 2023b). For the generation module, we compare the multi-agent frameworks using win rate - the ratio of times a generation framework is selected as the best by an LLM evaluator when comparing outputs from all generation workflows. After the extensive evaluation of both modules, we used the picked strategies to generate synthetic PO datasets. We set the temperature to 0 in all our evaluations to ensure reproducibility.
@@ 3.2 LLM-as-Evaluator
With the aim of automating the evaluation component of PO dataset generation, we assess the performance of LLMs in the role of evaluators using the Alpaca Eval, FairEval, PandaLM-Eval, and MT-Bench datasets. Our goal is to determine whether multi-agent workflows work better than a single agent for LLM evaluation. The system prompts for this task are modified version of the prompts used by Zheng et al. 2023a and are given in Appendix A.
@@
For each of these prompting strategy, we systematically analyze the performance of the LLMs by calculating Cohen’s Kappa, against the human annotations. The system prompts are given in Table 7 in Appendix A.
@@
We extend the evaluation from the LLM-as-a-Judge approach by forming juries composed of multiple LLMs. We test all possible permutations of the jury configurations. We use three datasets: FairEval, PandaLM-Eval and MT-Bench datasets for a more comprehensive analysis. We systematically analyze the performance of each jury configuration, focusing on how the size and combination of the LLMs affect their judgment accuracy. The Combined Scoring system prompt in Table 7 in Appendix A is used for all the jurors because it performed the best in our previous evaluation.
@@
We also evaluate the LLM Debate framework following the implementation described by Chan et al. 2023. In this approach, we assign three distinct roles—Psychologist, General Public, and Critic—and the three agents debate the scores that should be assigned to candidate responses. After the debate, each agent gives its final score which is used to determine which candidate response they vote for. These votes are then used to pick the best response. This strategy is evaluated using the FairEval, PandaLM-Eval, and MT-Bench benchmarks. Figure 1 illustrates the debate workflow employed in our study. The system prompt, the user message structure and the prompts for the roles used are given in Table 8 and Table 9 in Appendix A.
Figure 1: LLM Debate for evaluation
@@ 3.3 LLM-as-Generator
To evaluate the LLM Feedback Loop workflow for the generation module, we test different configurations using Llama-3.1-8b (Meta 2024) and Gemma-2-9b (Google 2024) models. In this framework, a generator LLM produces a response, which is then evaluated by a feedback LLM that provides improvement suggestions as shown in Figure 2. The generator revises the response based on these suggestions, and the process repeats for multiple iterations. The system prompt for the generator and reviewer is given in Table 10 and 11 in Appendix A. We calculate the win rate against single-agent GPT-4o (OpenAI 2024), Llama-3.1-8b and Gemma-2-9b baseline outputs on a subset of 500 prompts from the Argilla Capybara DPO dataset [^2] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized to identify the best configuration. We test the following configuration:
@@ 3.4 Preference Optimization Dataset
We use Llama-3.1-8b and Gemma-2-9b in the generation module and GPT-4o in the evaluation module to generate multiple DPO datasets and KTO datasets for single-agent improvement and multi-agent improvement. The prompts used for single-agent improvement dataset generation are given in Table 7, 10 and 11 in Appendix A. The prompt used for multi-agent improvement dataset generation is given in Table 12 in Appendix A. The evaluation code, all the evaluation outputs and the generated datasets are publicly available on GitHub [^3] https://github.com/ulrs0/MA-PO.
Single-Agent Improvement.
We use the prompts from Argilla Capybara DPO dataset. The Feedback Loop framework generates $N$ responses (where $N$ is the number of iterations). LLM-as-Evaluator picks the best response from the candidates to create the DPO and KTO dataset.
Multi-Agent Improvement.
We use the prompts and the human-generated responses from the No Robots dataset (Rajani et al. 2023). The evaluator is given the human response (reference response) and LLM response and is asked to generate feedback based on the reference response. The generated feedback and the human response are used to create the PO dataset. The goal of this dataset is to improve both response generation and response evaluation. The structure of the dataset is given in Figure 3.
Figure 3: Dataset structure for the mult-agent improvement PO dataset. The blue boxes represent the accepted responses and the yellow boxes represent the rejected responses.

### v4

@@ Abstract
This paper presents a novel methodology for generating synthetic Preference Optimization (PO) datasets using multi-agent workflows. We evaluate the effectiveness and potential of these workflows in automating and enhancing the dataset generation process. PO dataset generation requires two modules: (1) response evaluation, and (2) response generation. In the response evaluation module, the responses from Large Language Models (LLMs) are evaluated and ranked - a task typically carried out by human annotators that we automate using LLMs. We assess the response evaluation module in a 2 step process. In step 1, we assess LLMs as evaluators using three distinct prompting strategies. In step 2, we apply the winning prompting strategy to compare the performance of LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate. Our evaluation shows that GPT-4o-as-a-Judge is more consistent across all datasets. For the response generation module, we use the identified LLM evaluator configuration and compare different configurations of the LLM Feedback Loop. We use the win rate to determine the best multi-agent configuration for generation. Experimenting with various configurations, we find that the LLM Feedback Loop, with Llama as the generator and Gemma as the reviewer, achieves a notable 71.8% and 73.8% win rate over single-agent Llama and Gemma, respectively. After identifying the best configurations for both modules, we generate our PO datasets using the above pipeline.
The Fellowship of the LLMs: Multi-Agent Workflows for Synthetic Preference Optimization Dataset Generation
Samee Arif1and Sualeha Farid2and Abdul Hameed Azeemi1  and Awais Athar3  and Agha Ali Raza1 1Lahore University of Management Sciences, 2University of Michigan - Ann Arbor 3EMBL European Bioinformatics Institute {samee.arif, abdul.azeemi, agha.ali.raza}@lums.edu.pk sualeha@umich.edu, awais@ebi.ac.uk
@@ 1 Introduction
Large Language Models (LLMs) demonstrate a range of Natural Language Processing (NLP) capabilities, including text generation, question answering, and language understanding. However, LLMs can sometimes deviate from user instructions and exhibit unintended behaviors Tamkin et al. (2021). To mitigate this problem and align the LLM outputs more closely with human preferences, techniques like Reinforcement Learning from Human Feedback (RLHF) are used, which involves fine-tuning LLMs using the reward signal from human preferences Christiano et al. (2017). Improved methods like Direct Preference Optimization (DPO) Rafailov et al. (2024) eliminate the need for fitting the reward model and are more stable and performant. In DPO, the preference optimization dataset requires a pair of accepted and rejected responses for each prompt. The accepted response is one that better aligns with the desired human preferences. Other techniques like Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) require each response to indicate whether it is good or bad (i.e., as a binary classification task) instead of pairwise preferences.
@@
In the process of constructing the dataset of human preferences, the evaluation and ranking of the outputs generated by LLMs are typically done by human annotators, who assess these outputs based on various criteria such as instruction following, helpfulness, relevance, accuracy, depth, and creativity. The PO dataset generation process is divided into two modules: response evaluation and response generation. The response evaluation module involves assessing and ranking responses generated by LLMs, while the response generation module focuses on creating responses that align with the identified preferences. This manual process, while effective, is labor-intensive, time-consuming, inconsistent, and subject to human biases. In this work, we thus ask the question, Can we use LLM agents to automate and improve response evaluation and generation for constructing preference optimization (PO) datasets?.
@@
For the response evaluation step, we leverage LLMs as evaluators and compare several configurations including LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate to pick the best evaluation strategy. The selected response evaluation module is used to evaluate and identify the optimal response generation module. Previously, single-agents have been used to generate the responses for PO datasets; however, we use a multi-agent framework for response generation, which allows us to generate more refined, higher-quality responses. The multi-agent approach uses the collaboration between multiple LLMs, where one agent can provide suggestions for improvements, and the other can revise the response based on the feedback. This iterative process leads to a thorough refinement of the generated content, ensuring that the final output better aligns with human preferences and expectations.
@@
In this framework, the response generation module produces several possible responses, and the response evaluation module selects the best one from the list to create the PO dataset. We present multiple DPO and KTO datasets with the focus is on generating datasets to improve the performance of individual LLMs. The primary aim of the datasets is to enhance the performance and capabilities of individual LLMs by providing high-quality PO training data that better aligns with human judgment and expectations. Our contributions can be summarized as follows:
@@
    We generate synthetic PO datasets for LLM improvement by combining the best configuration for the evaluation and generation module.
@@ 2.1 Preference Optimization
Preference Optimization has emerged as a pivotal technique for aligning model outputs with human preferences. Rafailov et al. (2024) introduce DPO, a method that simplifies solving the standard RLHF problem by converting it into a classification task, enabling the extraction of the optimal policy in a straightforward way. Hong et al. (2024) introduce ORPO algorithm that combines the traditional supervised fine-tuning and preference alignment stages into a single process. The dataset for DPO and ORPO require annotated preference pairs, where each pair consists of two model outputs labeled according to which one better aligns with human preferences. Ethayarajh et al. (2024) introduce KTO, a cost-effective approach to align Large Language Models (LLMs) with human feedback, improving performance without the need for preference pairs. Argilla Distilabel (Álvaro Bartolomé Del Canto et al., 2024) uses LLM to judge between the responses of two models to create synthetic PO datasets. The datasets are available on Hugging Face [^1] https://huggingface.co/argilla. To our knowledge, no one has yet explored the use of Multi-Agent workflows for the generation of PO datasets.
@@ 2.2 Agent Frameworks
Recently, there has been a growing interest in using LLM multi-agent frameworks for different tasks. Zheng et al. (2023a) presents an evaluation of LLM-as-a-Judge on the MT-Bench (Zheng et al., 2023b) and Chatbot Arena (Li et al., 2024). Their results reveal that strong LLM judges like GPT-4 can match both controlled and crowd-sourced human preferences well, achieving over 80% agreement, the same level of agreement between humans. Additionally, they evaluate several variants of Llama and Vicuna on the dataset. They study the limitations of LLM-as-a-judge, including position, verbosity, and self-enhancement biases, as well as limited reasoning ability. Verga et al. (2024) explore the use of LLMs-as-a-Jury. Their approach, a Panel of LLM evaluators (PoLL), composed of a larger number of smaller models outperforms a single large judge. They also show that the PoLL approach exhibits less intra-model bias as compared to LLM-as-a-Judge. They use Command-R, GPT, Claude-3, and Mistral families for their study. Additionally, they compare two prompting strategies: (1) reference-based scoring where they provide the LLM with a reference answer, and (2) candidate answer and pair-wise scoring where they ask the LLM to pick the better response from the candidate responses. PoLL outperforms single-agents on KILT (Petroni et al., 2021) and Chatbot Arena.
@@
Liang et al. (2024) introduce Multi-Agent Debate (MAD) to encourage divergent thinking in LLMs. They mitigate the Degeneration-of-Thought (DoT) problem, which is that once the LLM has established confidence in its solutions, it is unable to generate novel thoughts. In their approach, the affirmative LLM and the negative LLM debate on the answer while the LLM judge evaluates both arguments after each round of debate. They evaluate the approach on the Commonsense Machine Translation Dataset (Chinese to English) (He et al., 2020) and their Counter-Intuitive Arithmetic Reasoning (CIAR) dataset. MAD was able to achieve a 37% accuracy on the CIAR dataset using GPT-3.5-Turbo which outperforms Chain-of-Thought, Self-Consistency, and Self-Reflection prompting. They also show that using the MAD approach decreases bias and increases response diversity. Du et al. (2023) evaluates a different variant of multi-agent debate where multiple models generate their own responses, and each model receives the opinions of the other models, then updates its response if necessary. This is done for multiple rounds. Du et al. (2023) evaluates the approach on the following tasks: Biography generation, MMLU, Chess move validity and optimality, Arithmetic, and Grade school math,. Their approach using ChatGPT and Bard outperforms single-agent on all the tasks. To evaluate LLM responses Chan et al. (2023) presents another variant of multi-agent debate. Their architecture involves assigning agents different roles such as General Public, Critic, Psychologist, News Author, and Scientist. They used ChatGPT and GPT-4 for their evaluation on FairEval (Wang et al., 2023a) dataset and achieved a Cohen’s Kappa score of 0.40 using LLM Debate, 0.03 more than the single-agent.
@@ 3.1 Experimental Setup
In this study, we perform experiments on the three categories of models given in Table 1. For the evaluation module, we evaluate single-agents and multi-agent frameworks on four datasets, Alpaca Eval (Li et al., 2023), FairEval (Wang et al., 2023a), PandaLM-Eval (Wang et al., 2024, 2023b) and MT-Bench (Zheng et al., 2023b). For the generation module, we compare the multi-agent frameworks using win rate - the ratio of times a generation framework is selected as the best by an LLM evaluator when comparing outputs from all generation workflows. After the extensive evaluation of both modules, we used the picked strategies to generate synthetic PO datasets. We set the temperature to 0 in all our evaluations to ensure reproducibility.
@@ 3.2 LLM-as-Evaluator
With the aim of automating the evaluation component of PO dataset generation, we assess the performance of LLMs in the role of evaluators using the Alpaca Eval, FairEval, PandaLM-Eval, and MT-Bench datasets. Our goal is to determine whether multi-agent workflows work better than a single-agent for LLM evaluation. The system prompts for this task are modified version of the prompts used by Zheng et al. (2023a) and are given in Appendix A.
@@
For each of these prompting strategy, we systematically analyze the performance of the LLMs by calculating Cohen’s Kappa, against the human annotations. The system prompts are given in Table LABEL:tab:prompt-llm-judge in Appendix A. ko
@@
We extend the evaluation from the LLM-as-a-Judge approach by forming juries composed of multiple LLMs. Specifically, we test all possible combinations of the six LLM models when forming juries of sizes ranging from 2 to 6. We use three datasets: FairEval, PandaLM-Eval and MT-Bench datasets for a more comprehensive analysis. We systematically analyze the performance of each jury configuration, focusing on how the size and combination of the LLMs affect their judgment accuracy. The Combined Scoring system prompt in Table LABEL:tab:prompt-llm-judge in Appendix A is used for all the jurors because it performed the best in our previous evaluation.
Figure 1: LLM Debate for evaluation
@@
We also evaluate the LLM Debate framework following the implementation described by Chan et al. (2023). In this approach, we assign three distinct roles—Psychologist, General Public, and Critic—and the three agents debate the scores that should be assigned to candidate responses. After the debate, each agent gives its final score which is used to determine which candidate response they vote for. These votes are then used to pick the best response. This strategy is evaluated using the FairEval, PandaLM-Eval, and MT-Bench benchmarks. Figure 1 illustrates the debate workflow employed in our study. The system prompt, the user message structure and the prompts for the roles used are given in Table LABEL:tab:prompt-llm-debate and Table LABEL:tab:prompt-roles in Appendix A.
@@ 3.3 LLM-as-Generator
To evaluate the LLM Feedback Loop workflow for the generation module, we test different configurations using Llama-3.1-8b (Meta, 2024) and Gemma-2-9b (Google, 2024) models. In this framework, a generator LLM produces a response, which is then evaluated by a feedback LLM that provides improvement suggestions as shown in Figure 2. The generator revises the response based on these suggestions, and the process repeats for multiple iterations. The system prompt for the generator and reviewer is given in Table LABEL:tab:generator_prompt and LABEL:tab:reviewer_prompt in Appendix A. We calculate the win rate against single-agent GPT-4o (OpenAI, 2024), Llama-3.1-8b and Gemma-2-9b baseline outputs on a subset of 500 prompts from the Argilla Capybara DPO dataset [^2] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized to identify the best configuration. We test the following configuration:
@@ 3.4 Preference Optimization Dataset
We use the best configurations of the generation and evaluation modules to generate the DPO and KTO datasets. The generation module produces $N$ responses (where $N$ is the number of feedback iterations), which are then passed to the evaluation module. The evaluation module sorts these responses into the accepted and rejected fields in the DPO and KTO datasets. In this study, we use the prompts from the Argilla Capybara DPO dataset. The prompt templates used for LLM improvement dataset generation are given in Table LABEL:tab:prompt-llm-judge, LABEL:tab:generator_prompt and LABEL:tab:reviewer_prompt in Appendix A. The evaluation code, all the evaluation outputs and the generated datasets are publicly available on GitHub [^3] https://github.com/ulrs0/MA-PO.